# General Hamiltonian description of nonreciprocal interactions

### v3

@@ Abstract
In a vast class of systems, which includes members as diverse as active colloids and bird flocks, interactions do not stem from a potential, and are in general nonreciprocal. Thus, it is not possible to define a conventional energy function, nor to use analytical or numerical tools that rely on it. Here, we overcome these limitations by constructing a Hamiltonian that includes auxiliary degrees of freedom; when subject to a constraint, this Hamiltonian yields the original non-reciprocal dynamics. We show that Glauber dynamics based on the constrained Hamiltonian reproduces the steady states of the original Langevin dynamics, as we explicitly illustrate for dissipative XY spins with vision-cone interactions. Further, the symplectic structure inherent to our construction allows us to apply the well-developed notions of Hamiltonian engineering, which we demonstrate by varying the amplitude of a periodic drive to tune the spin interactions between those of a square and a chain lattice geometry. Overall, our framework for generic nonreciprocal pairwise interactions paves the way for bringing to bear the full conceptual and methodological power of conventional statistical mechanics and Hamiltonian dynamics to nonreciprocal systems.
@@ I Introduction
Newton’s third law states that every action has a reaction equal in magnitude and opposite in direction. For two particles interacting via a potential, the interaction forces are reciprocal: $\mathbf{F}_{2\to 1}=-\mathbf{F}_{1\to 2}$. However, for many mesoscopic and macroscopic particles, interactions do not arise from a potential but are instead mediated by a non-equilibrium field [1]. For example, interactions between colloids can be mediated by hydrodynamic flows, by a dissolved chemical, or by a temperature field [2, 3, 4, 5, 6, 7]. Respectively, interactions between animals, such as birds, wildebeest or humans, can be mediated by their vision field [8, 9, 10]. In such systems, interactions are in general nonreciprocal: $\mathbf{F}_{2\to 1}\neq-\mathbf{F}_{1\to 2}$. Nonreciprocal interactions have been reported in systems including spinning starfish embryos [11] as well as active colloids and droplets [12, 13, 14, 15, 16, 17, 18, 19, 20]; they have also been programmed in colloids [21, 22] and in robots [23, 24, 25]. Despite recent advances [26, 27, 25, 28, 29, 30, 31], a general statistical-mechanical framework for nonreciprocal systems is still missing [26].
[…]
Figure 1: Classical systems can be classified as either Hamiltonian (e.g., harmonic oscillator) or non-Hamiltonian. The latter obey forces which can either be derived from a potential (e.g., sedimenting particle) or not (e.g., nonreciprocal systems). Here we construct a Hamiltonian embedding for nonreciprocal systems.
@@
Non-Hamiltonian systems, on the other hand, lack conjugate variables, and hence do not enjoy any of the advantages offered by the symplectic structure. Broadly speaking, they fall into two categories: systems whose equations of motion are governed by conservative forces that can be derived from a potential, and those that cannot, cf. Fig. 1. Perhaps the most common example of a non-Hamiltonian potential system is a particle subject to a conservative force, such as gravity, and coupled to a dissipative medium, such as a liquid  [^1] We define dissipation by lack of conservation of phase-space volume under the dynamics.. A notable advantage of having a potential is the ability to use Monte-Carlo methods [38] to investigate the properties of the system coupled to a thermal bath.
@@
Conversely, in the simplest setting, nonreciprocal systems are modeled by forces that cannot be derived from a single mutual interaction potential [26]. These nonconservative forces do not admit an energy function and are, furthermore, non-Hamiltonian in general. Physically, this means that nonreciprocal systems are intrinsically out of thermal equilibrium, which significantly complicates their study beyond the bare simulation of the underlying equations of motion. For instance, due to the absence of a potential, it is a priori unclear how to apply Monte-Carlo methods to simulate and analyze properties of their non-equilibrium steady states. Similarly, due to the lack of conjugate variables, deriving effective descriptions for nonreciprocal models cannot be done using canonical transformations.
@@
In this work, we remove these obstacles by designing a Hamiltonian embedding for nonreciprocal systems. Specifically, for any equation of motion with pairwise interactions described by nonreciprocal velocity-independent forces, we present a construction of a reciprocal Hamiltonian, comprised of the original system and an auxiliary degree of freedom for each original one [Fig. 2]. By imposing a constraint on each (original, auxiliary)-pair, the equation of motion derived from this Hamiltonian reproduces the original nonreciprocal dynamics. Our results apply equally to inertial and dissipative nonreciprocal dynamics; for dissipative nonreciprocal systems in particular, the original and auxiliary degrees of freedom form canonically conjugate variables.
@@
We exemplify our construction using dissipative XY spins on a square lattice with vision-cone interactions — a paradigmatic example of nonreciprocal interactions [39, 29, 40, 41]. To demonstrate the utility of our construction, we first show (both analytically and numerically) the equivalence of the steady states obtained from Monte-Carlo simulations with Glauber dynamics based on the Hamiltonian embedding on the one hand, and the original equations of motion subject to Langevin noise on the other. We then use the emergent symplectic structure of the Hamiltonian embedding to analyze the vision-cone XY system subject to a periodic drive using Floquet theory [42], and derive effective equations of motion for the driven system. In the SM [43], we provide a general proof for the construction, accompanied by applications to four experimental systems featuring nonreciprocal interactions: metal-dielectric Janus particles [15, 20], walker robots [44], robotic metamaterials [23], and feedback-controlled active particles [21].
@@ II XY spins with nonreciprocal Vision-Cone Interactions
where the Boltzmann constant $k_{B}{=}1$, and the damping coefficient $\gamma{=}1$. The corresponding vision-cone XY interaction [Fig. 2(a)] gives rise to a minimal model for nonreciprocal interactions, which has recently gained significant attention. Previous work has studied the role of non-reciprocity on the transition from the disordered to the ferromagnetic phase [39, 29, 40, 41], as well as on the creation and annihilation of topological defects [45, 40].
@@ III Constrained Hamiltonian Embedding
[…]
@@
The second term, $H_{Sa}$, represents the interaction between the two layers, where the auxiliary spins $\mathbf{a}_{i}$ interact with the original spins $\mathbf{S}_{j}$, following
@@
The magnitude of each of the two terms in $H_{Sa}$ above directly corresponds to that in the symmetrized interaction, see Fig. 2(b) inset. As we show next, this is essential to recover the original nonreciprocal dynamics.
@@
it is preserved under the equations of motion, i.e., $\theta_{i}(t){-}\varphi_{i}(t){=}\pi$ for all times, as we show in the SI [43]. Because, under the constraint, the dynamics of the $\varphi_{i}(t)$ degree of freedom is tied to that of the $\theta_{i}(t)$ spins, the equations of motion (4) uncouple. Moreover, each degree of freedom recovers precisely the nonreciprocal vision-cone interaction term from Eq. (1). Formally, the constraint (5) causes the unwanted terms in the last two lines of Eq. (4) to cancel, and makes the remaining equations the same.
Note that the form of the Hamiltonian embedding is not unique. For instance, one can distribute the interaction symmetrically among the two types of degrees of freedom without changing the resulting equations of motion upon imposing the constraint. It is currently an open question whether this freedom can be used to imprint additional properties into the embedding Hamiltonian.
@@
The procedure illustrated above is not limited to vision-cone XY interactions. In the SI [43], we first present a general proof for arbitrary velocity-independent nonreciprocal interactions. To illustrate the wide applicability of our framework, we construct the Hamiltonian for four specific examples: metal-dielectric Janus particles [20, 15], self-aligning active particles such as walker robots [44, 46], robotic metamaterials [23], and feedback-controlled active particles [21]. In these examples, the dynamics are overdamped, i.e., dissipation dominates over inertia. However, using suitable kinetic-energy terms, the Hamiltonian embedding also applies to inertial (i.e., non-dissipative) nonreciprocal dynamics, suggesting a broad applicability, see SI [43].
@@
In the following, we illustrate the utility of the proposed Hamiltonian description by (i) demonstrating that the long-time steady state of the original Langevin dynamics (1) coincides with the steady state of Glauber dynamics generated by the Hamiltonian (2), thus enabling the use of Monte-Carlo methods to study nonreciprocal dynamics. In addition, (ii) we engineer the strength of vision cone interactions along different lattice directions with the help of high-frequency periodic drives, explicitly using the symplectic structure of the Hamiltonian formulation; this enables the application of Floquet engineering to control the phase transitions of nonreciprocal systems.
@@
IV Equivalence between Hamiltonian Monte-Carlo and Langevin Dynamics
@@
The absence of a well-defined interaction energy per bond precludes the direct application of Monte-Carlo methods to study nonreciprocal systems. To circumvent this problem, recent works have introduced new concepts and constructions [47, 48, 49] to perform Monte-Carlo simulations, such as the ‘selfish energy’ of a lattice site $i$, $E_{i}{=}{-}\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i})\cos(\theta_{i}{-}\theta_{j})$ [30, 29, 50], or modified acceptance rates that include nonreciprocal interactions through a Legendre-like transform similar to a grand-canonical ensemble [51]. A notable common disadvantage is that these tools are ad-hoc, i.e., it is a priori unclear whether the resulting steady states correspond to those of the physical microscopic Langevin dynamics. Additionally, they do not enjoy formal mathematical guarantees for convergence, unlike Monte-Carlo methods for reciprocal systems [38].
@@
Starting from the reciprocal Hamiltonian in Eq. (2), we propose a two-step Monte-Carlo algorithm. First, we apply Glauber dynamics at temperature $T$ defined by the transition rates [52]
$w\left(\theta_{i}\rightarrow\theta_{i}^{\prime}\right)=\frac{1}{2}\left(1-\tanh\frac{H(\theta_{i}^{\prime},\varphi_{i})-H(\theta_{i},\varphi_{i})}{2T}\right),$ (6)
to the original $\mathbf{S}$-spin layer while keeping the auxiliary spin degree of freedom $\varphi_{j}$ fixed. In a second step, we enforce the constraint by deterministically adjusting the auxiliary spins according to the accepted move: $\varphi_{i}{\rightarrow}\varphi^{\prime}_{i}{=}\theta_{i}^{\prime}{+}\pi$; this causes the sampler to converge to a non-equilibrium steady state. Our complete algorithm is provided in [43]. Note that sampling under the constraint is similar to directed loop algorithms, which are used to explore the extensive discrete set of ground states in spin models supporting an emergent gauge theory [53], and more efficient algorithms may exist.
The long-time state of the above Glauber dynamics is imposed by the steady-state condition
$\Big{\{}\rho(\theta_{i}+\delta,\varphi_{i}+\delta)w(\theta_{i}+\delta\rightarrow\theta_{i})$
$-\rho(\theta_{i},\varphi_{i})w(\theta_{i}\rightarrow\theta_{i}+\delta)\Big{\}}\Big{|}_{\text{constr.}}=0$ (7)
where $\delta{=}\theta^{\prime}_{i}{-}\theta_{i}$ denotes the small change of the angle $\theta_{i}$ in the Glauber update. We note that detailed balance here is explicitly broken when restoring the constraint, see SM [43]. Using the rates from Eq. (6) and Taylor-expanding to linear order in $\delta$, the density $\rho$ obeys
$\left(\partial_{\varphi_{i}}{+}\partial_{\theta_{i}}\right)\left[\partial_{\theta_{i}}H|_{\text{constr.}}{+}T\left(\partial_{\varphi_{i}}{+}\partial_{\theta_{i}}\right)\right]\rho\left(\theta_{i},\varphi_{i}\right){=}0.$ (8)
We now argue that Eq. (8) arises independently as the Fokker-Planck equation of the original Langevin dynamics at steady state. Recall that the equation of motion (1) is generated by the Hamiltonian (2) under the constraint (5) for each pair $(\theta_{i},\varphi_{i})$. In addition, assume that the two spins $\mathbf{S}_{i}$ and $\mathbf{a}_{i}$ are coupled to the same thermal bath and therefore experience identical noise. Let us rewrite the overdamped Langevin dynamics (1) as $\dot{\alpha}_{i}\left(t\right){=}{-}\left\{\alpha_{i},H\right\}+\sqrt{2T}\eta_{i}(t)$ under the constraint, for $\alpha{=}\theta,\varphi$. One way to recover the configuration-space distribution is by introducing a test function $f(\{\alpha_{j}(t)\})$ and calculating its average value under an ensemble of trajectories generated by the above Langevin equation. In a steady state, the ensemble-average of $f$ is time-independent, i.e., $\partial_{t}\left\langle f(\{\alpha_{j}(t)\})\right\rangle{=}0$. Denoting the corresponding phase-space probability density by $\rho\left(\theta_{i},\varphi_{i}\right)$, this time-translation invariance leads to the textbook steady-state Fokker-Planck equation when the constraint is ignored [54, 55]. However, taking into account the constraint (5) imprints additional correlations between $\theta_{j}$ and $\varphi_{j}$, leading to Eq. (8), see SI [43]. Thus, Glauber Monte-Carlo dynamics generated by the Hamiltonian (2) produces the same steady-state distribution as the Langevin equation (1).
@@
Figure 3: (a) Average magnetization $\langle m\rangle$ and (b) specific heat $C_{v}$ as functions of temperature $T$, obtained from Monte-Carlo simulations with Glauber dynamics based on either the selfish energy (red) or the constrained Hamiltonian embedding (blue), as well as simulations of the original Langevin dynamics (green). Whereas the constrained Glauber and Langevin simulations show the same critical temperature within the statistical uncertainty: $T_{c}=0.78\pm 0.02J$ (Langevin) and $T_{c}=0.79\pm 0.03J$ (constrained Glauber), the selfish-energy Glauber dynamics predicts $T_{c}=0.86\pm 0.02J$. The square lattice has linear dimension $L{=}100$. We use a vision cone angle $\Psi{=}0.85\pi$ and set the time step $\delta t=0.1$ for the Langevin simulation. We use $20L^{4}$ ($3\times 10^{5}$) timesteps for the constrained Glauber (Langevin) dynamics.

### v4

# Hamiltonian description of nonreciprocal interactions

@@ Abstract
In a vast class of systems, which includes members as diverse as sedimenting particles and bird flocks, interactions do not stem from a potential, and are in general nonreciprocal. Thus, it is not possible to define a conventional energy function, nor to use analytical or numerical tools that rely on it. Here, we overcome these limitations by constructing a Hamiltonian that includes auxiliary degrees of freedom; when subject to a constraint, this Hamiltonian yields the original nonreciprocal dynamics. We show that Glauber dynamics based on the constrained Hamiltonian reproduce both stationary and nonstationary states of the original Langevin dynamics, as we explicitly illustrate for dissipative XY spins with vision-cone interactions. Further, the symplectic structure inherent to our construction enables us to apply the well-developed notions of Hamiltonian engineering, which we demonstrate by varying the amplitude of a periodic drive to tune the spin interactions between those of a square and a chain lattice geometry. Overall, our framework for generic nonreciprocal pairwise interactions paves the way for bringing to bear the full conceptual and methodological power of conventional statistical mechanics and Hamiltonian dynamics to nonreciprocal systems.
@@ I Introduction
Newton’s third law states that every action has a reaction equal in magnitude and opposite in direction. For two particles interacting via a potential, the interaction forces are reciprocal: $\mathbf{F}_{2\to 1}=-\mathbf{F}_{1\to 2}$. However, for many mesoscopic and macroscopic particles, interactions do not arise from a potential but are instead mediated by a nonequilibrium field [16, 31]. For example, interactions between colloids can be mediated by hydrodynamic flows, by a dissolved chemical, or by a temperature field [28, 37, 77, 78, 49, 2]. Respectively, interactions between animals, such as birds, wildebeest or humans, can be mediated by their vision field [66, 9, 27]. In such systems, interactions are in general nonreciprocal: $\mathbf{F}_{2\to 1}\neq-\mathbf{F}_{1\to 2}$. Nonreciprocal interactions have been reported in systems including sedimenting particles [64, 31, 75, 74], spinning starfish embryos [82] as well as active colloids and droplets [69, 76, 56, 89, 86, 39, 13, 53, 21]; they have also been programmed in colloids [11, 48] and in robots [17, 34, 30]. Despite recent advances [44, 46, 30, 23, 51, 5, 43], a general statistical-mechanics framework for nonreciprocal systems is still missing [44].
To appreciate the reason for this, note that in classical and statistical mechanics, one may distinguish between Hamiltonian and non-Hamiltonian systems, see Fig. 1. Hamiltonian systems possess conjugate pairs of variables that span phase space (e.g., position and momentum). Mathematically, this is expressed by the Poisson bracket $\{\cdot,\cdot\}$ which endows phase space with a so-called symplectic structure [4]. This enables the use of canonical transformations [19, 14], which preserve the Poisson bracket, to find integrals of motion or derive simplified effective models that approximate the dynamics [14, 40, 70].
Figure 1: Classical systems can be classified as either Hamiltonian (e.g., two electric charges in vacuum, left panel) or non-Hamiltonian. The latter can have interactions that may be derived from a potential (e.g., two electric charges on a frictional substrate, middle panel) or not (e.g., birds with vision-cone interactions, right panel). In this work, we construct a Hamiltonian embedding for nonreciprocal systems.
@@
Non-Hamiltonian systems, on the other hand, lack conjugate variables, and hence do not enjoy any of the advantages offered by the symplectic structure. Broadly speaking, they fall into two categories: systems whose equations of motion are governed by conservative forces that can be derived from a potential, and those that cannot, cf. Fig. 1. A simple example of a non-Hamiltonian system with interactions arising from a potential is a pair of particles subject to dissipation, for example through friction with a substrate (Fig. 1, middle panel)  [^1] We define dissipation by lack of conservation of phase-space volume under the dynamics.. A notable advantage of having a potential is the ability to use Monte-Carlo methods [47] to investigate the properties of the system coupled to a thermal bath.
@@
Conversely, in the simplest setting, nonreciprocal systems are modeled by forces that cannot be derived from a single mutual interaction potential [44]. These nonconservative forces do not admit an energy function and are, furthermore, non-Hamiltonian in general. Physically, this means that nonreciprocal systems are intrinsically out of thermal equilibrium, which significantly complicates their study beyond the bare simulation of the underlying equations of motion. For instance, due to the absence of a potential, it is a priori unclear how to apply Monte-Carlo methods to simulate and analyze properties of their nonequilibrium steady states. Similarly, due to the lack of conjugate variables, deriving effective descriptions for nonreciprocal models cannot be done using canonical transformations.
@@
In this work, we remove these obstacles by designing a Hamiltonian embedding for nonreciprocal systems. Specifically, for any equation of motion with pairwise interactions described by nonreciprocal forces, we present a construction of a reciprocal Hamiltonian, comprised of the original system and an auxiliary degree of freedom for each original one [Fig. 2]. By imposing a constraint on each (original, auxiliary)-pair, the equation of motion derived from this Hamiltonian reproduces the original nonreciprocal dynamics. Our results apply equally to inertial and dissipative nonreciprocal dynamics; for dissipative nonreciprocal systems in particular, the original and auxiliary degrees of freedom form canonically conjugate variables.
@@
We exemplify our construction using dissipative XY spins on a square lattice with vision-cone and chase-and-run interactions – paradigmatic examples of nonreciprocal interactions [20, 51, 62, 25]. To demonstrate the utility of our construction, we first show (both analytically and numerically) the equivalence of the states, either steady or nonstationary, obtained from Monte-Carlo simulations with Glauber dynamics based on the Hamiltonian embedding on the one hand, and the original equations of motion subject to Langevin noise on the other. We then use the emergent symplectic structure of the Hamiltonian embedding to analyze the vision-cone XY system subject to a periodic drive using Floquet theory [18], and derive effective equations of motion for the driven system. In the SM [1], we provide a general proof for the construction, accompanied by applications to five experimental systems featuring nonreciprocal interactions: metal-dielectric Janus particles [89, 21], walker robots [6], robotic metamaterials [17], feedback-controlled active particles [11], and sedimenting particles [64].
@@ II XY spins with nonreciprocal Vision-Cone Interactions
where the Boltzmann constant $k_{B}{=}1$, and the damping coefficient $\gamma{=}1$. The corresponding vision-cone XY interaction [Fig. 2(a)] gives rise to a minimal model for nonreciprocal interactions, which has recently gained significant attention. Previous work has studied the role of nonreciprocity on the transition from the disordered to the ferromagnetic phase [20, 51, 62, 25], as well as on the creation and annihilation of topological defects [67, 62].
@@ III Constrained Hamiltonian Embedding
with conjugate degrees of freedom $\{\theta_{i},\varphi_{j}\}{=}\delta_{ij}$, where $\{\cdot,\cdot\}$ is the Poisson bracket; hence, $\varphi_{i}$ acts as the canonical momentum for $\theta_{i}$, see SI [1] for details.
@@
The second term, $H_{Sa}$, represents the coupling between the two layers, where the auxiliary spins $\mathbf{a}_{i}$ interact with the original spins $\mathbf{S}_{j}$, following
@@
The magnitude of each of the two terms in $H_{Sa}$ above directly corresponds to that in the symmetrized interaction, see Fig. 2(b) inset; as we show next, this is essential to recover the original nonreciprocal dynamics.
@@
it is preserved under the equations of motion, i.e., $\theta_{i}(t){-}\varphi_{i}(t){=}\pi$ for all times, as we show in the SI [1]. Because, under the constraint, the dynamics of the $\varphi_{i}(t)$ degree of freedom is tied to that of the $\theta_{i}(t)$ spins, the equations of motion (4) uncouple. Moreover, each degree of freedom recovers precisely the nonreciprocal vision-cone interaction term from Eq. (1). Formally, the constraint (5) causes the unwanted terms in the last two lines of Eq. (4) to cancel, and makes the remaining equations identical  [^2] Notice that, after applying the constraint, the vision cones for $\varphi_{i}$ and $\theta_{i}$ point in the same directions, even though $\theta_{i}-\varphi_{i}=\pi$..
@@
The procedure illustrated above is not limited to vision-cone XY interactions. In the SI [1], we first present a general proof for arbitrary nonreciprocal interactions. To illustrate the wide applicability of our framework, we construct the Hamiltonian for five specific examples: metal-dielectric Janus particles [21, 89], self-aligning active particles such as walker robots [6, 7], robotic metamaterials [17], feedback-controlled active particles [11], and sedimenting particles [64]. In these examples, the dynamics are overdamped, i.e., dissipation dominates over inertia. However, using suitable kinetic-energy terms, the Hamiltonian embedding also applies to inertial (i.e., non-dissipative) nonreciprocal dynamics, suggesting a broad applicability, see SI [1].
According to Liouville’s theorem, without constraint, the Hamiltonian embedding gives rise to nondissipative dynamics in the embedding phase space, even if the original equations of motion are dissipative. This feature is brought about by the symplectic structure inherent to the Hamiltonian embedding (e.g., in the vision-cone XY spins that we discussed, $\theta$ and $\varphi$ are conjugate variables). By contrast, on the submanifold defined by the constraint, the dynamics are dissipative. This apparent contradiction is resolved by noticing that the constrained submanifold is a set of measure zero in the embedding phase space which does not contribute to phase-space volume.
As a direct consequence of the constraint, trajectories with initial conditions on the constraint manifold remain confined to it for all time, and hence never intersect with trajectories that start outside. Therefore, dynamics outside the manifold, which are reciprocal, do not affect the dynamics on the constraint manifold, which are nonreciprocal. Thence, the phase-space probability distribution evolves independently on the constraint manifold and outside it. It is therefore possible for an initial ensemble on the constraint manifold to evade thermalization, even when the embedding Hamiltonian dynamics outside the manifold are ergodic. This provides a way to embed nonequilibrium configuration-space distributions of nonreciprocal systems into the larger Hamiltonian phase space. Because the nonreciprocal manifold is a set of measure zero within the total phase space, it does not affect expectation values of observables defined on the total phase space. Hence, observables on the embedding phase space are insensitive to non-stationary behaviors on the nonreciprocal manifold, such as the breaking of continuous time-translation symmetry or persistent configuration-space currents.
Furthermore, notice that, when evaluated on the constraint (5), the total Hamiltonian $H$ vanishes identically; therefore, $H$ does not describe the energy of the original nonreciprocal system, but merely provides a generator for its dynamics. Moreover, the ground-state manifold of $H$ need not obey the constraint. These unusual features reflect the inherent out-of-equilibrium character of the original nonreciprocal system.
Finally, we note that the form of the Hamiltonian embedding is not unique. For instance, one can distribute the interaction symmetrically among the two types of degrees of freedom without changing the resulting equations of motion upon imposing the constraint. It is currently an open question whether this freedom can be used to imprint additional properties or constraints into the embedding Hamiltonian.
@@
In what follows, we illustrate the utility of the proposed Hamiltonian description by (i) demonstrating that the long-time steady nonstationary state of the original Langevin dynamics (1) coincides with the steady nonstationary state of Glauber dynamics generated by the Hamiltonian (2), thus enabling the use of Monte-Carlo methods to study nonreciprocal dynamics. In addition, (ii) we engineer the strength of vision cone interactions along different lattice directions with the help of high-frequency periodic drives, explicitly using the symplectic structure of the Hamiltonian formulation; this enables the application of Floquet engineering to control the phase transitions of nonreciprocal systems.
@@
IV Equivalence between Hamiltonian Monte-Carlo and Langevin Dynamics Subject to White Noise
@@
The absence of a well-defined interaction energy per bond precludes the direct application of Monte-Carlo methods to study nonreciprocal systems. To circumvent this problem, recent works have introduced new concepts and constructions [85, 63, 71] to perform Monte-Carlo simulations, such as the ‘selfish energy’ of a spin on the lattice site $i$, $E_{i}{=}{-}\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i})\cos(\theta_{i}{-}\theta_{j})$ [5, 51, 15], or modified acceptance rates that include nonreciprocal interactions through a Legendre-like transform similar to a grand-canonical ensemble [60]. A notable common disadvantage of these tools is that it is a priori unclear whether the resulting steady states correspond to those of the microscopic Langevin dynamics; moreover, many nonreciprocal systems evolve into nonstationary states. Additionally, these methods do not enjoy formal mathematical guarantees for convergence, unlike Monte-Carlo methods for reciprocal systems [47].
@@
Starting from the reciprocal Hamiltonian embedding, we propose a Monte-Carlo algorithm to sample nonequilibrium steady states and nonstationary states of nonreciprocal systems. Specifically, we apply constrained Glauber dynamics [35] at temperature $T$ defined by the transition rates $w\left(\theta_{i}\rightarrow\theta_{i}^{\prime}\right)={1}/{2}\left[1-\tanh(\Delta E(\theta_{i},\theta^{\prime}_{i})/(2T))\right]$ with
$\Delta E(\theta_{i},\theta^{\prime}_{i})=\frac{1}{2}\big\{H(\theta^{\prime},\varphi)-H(\theta,\varphi^{\prime})\big\}\big|_{\text{constr.}}\,.$ (6)
Here, the notation $\{\cdot\}|_{\text{constr.}}$ means that the operations in the bracket are carried out first, and only then is the constraint applied. As a result, the auxiliary variable $\varphi_{i}$ drops out of the expression for the energy difference $\Delta E(\theta_{i},\theta^{\prime}_{i})$. Consequently, we only need to simulate the dynamics of the original $\mathbf{S}$-spin layer, thus keeping the number of simulated degrees of freedom equal to that of the original system.
For large system sizes, $L{\gg}1$, these Glauber dynamics give rise to a Fokker-Planck equation for the configuration-space distribution $\rho(\vec{\theta},t_{\text{MC}})$ (see Methods):
$\frac{\partial\rho(\vec{\theta},t_{\textrm{MC}})}{\partial t_{\textrm{MC}}}=\alpha\sum_{i=1}^{L^{2}}\partial_{\theta_{i}}\Big[T\partial_{\theta_{i}}\rho(\vec{\theta},t_{\textrm{MC}}){-}F_{i}(\vec{\theta})\rho(\vec{\theta},t_{\textrm{MC}})\Big],$
where $F_{i}(\vec{\theta}){=}{-}\partial_{\theta_{i}}H(\vec{\theta},\vec{\varphi})\big|_{\text{constr.}}$ is the (nonreciprocal) force acting on spin $\mathbf{S}_{i}$. Here $\alpha$ is a scaling factor relating the physical time and Monte-Carlo time: $t=\alpha t_{\text{MC}}$, set by temperature $T$ and the support size $\Delta$ of the uniform distribution according to which the Glauber updates are performed (see Methods).
In the Methods section, we prove that Eq. (IV) coincides with the Fokker-Planck equation obtained from the original Langevin equations of motion. Therefore, the Hamiltonian embedding allows us to define transition rates that enable the use of the Glauber algorithm in nonreciprocal systems. Our Monte-Carlo scheme allows us to sample nonequilibrium states since detailed balance is broken by the constraint, see Eq. (6) and SM [1]. Hence, we can bypass brute-force Langevin dynamics and directly sample both nonequilibrium steady states as well as oscillatory nonstationary states, as we now demonstrate in numerical simulations.
@@
Figure 3: Nonequilibrium steady state in XY spins with vision-cone interactions. (a) Average magnetization $\langle m\rangle$ and (b) specific heat $C_{v}$ as functions of temperature $T$, obtained from Monte-Carlo simulations with Glauber dynamics based on either the selfish energy (green) or the constrained Hamiltonian embedding (blue), as well as simulations of the original Langevin dynamics (red). Whereas the constrained Glauber and Langevin simulations show the same critical temperature within the statistical uncertainty: $T_{c}/J=0.79\pm 0.04$ (Langevin) and $T_{c}/J=0.79\pm 0.04$ (constrained Glauber), the selfish-energy Glauber dynamics predicts $T_{c}/J=0.86\pm 0.02J$. The square lattice has linear dimension L=100L{=}100, the vision cone angle is $\Psi{=}0.85\pi$, and averages are performed over $\mathcal{N}{=}100$ trajectories. For the Langevin dynamics, we set a time step $J\delta t{=}0.01$ and use $100L^{2}$ timesteps. In the constrained Glauber algorithm[1], we set a support size $\Delta{=}0.1$, and use use $200L^{4}$ timesteps.
IV.1 Nonequilibrium steady states
For the vision-cone spins from Eq. (1), using the Hamiltonian embedding (2) we compute the energy difference entering the constrained Glauber transition rates as $\Delta E(\theta_{i},\theta^{\prime}_{i}){=}{-}\frac{1}{2}\sum_{j\in\langle ij\rangle}[J_{ij}(\theta_{i}^{\prime}){+}J_{ij}(\theta_{i})]\times\qquad\qquad[\cos(\theta_{i}^{\prime}{-}\theta_{j}){-}\cos(\theta_{i}{-}\theta_{j})]$.